# Dispersion of first sound in a weakly interacting ultracold Fermi liquid: an exact calculation

### v2

@@ Abstract
In this regime, collisions bring the quasiparticle distribution back to equilibrium much faster than the typical time $1/\omega_{0}$ at which the sound wave evolves. Only the few components of the distribution that are not affected by collisions (i.e. those that belong to the zero-energy space of the collision kernel) remain significantly excited.
To obtain the exact solution of the transport equation in this limit, we expand the distribution $\nu$ over a basis of orthogonal polynomials
$\nu(\epsilon,\theta)=\sum_{n,l}\nu_{n}^{l}P_{l}(\cos\theta)Q_{n}(\epsilon)$ (13)
@@
$\int_{-\infty}^{+\infty}g(\epsilon)Q_{n}(\epsilon)Q_{m}(\epsilon)\textrm{d}\epsilon=||Q_{n}||^{2}\delta_{nm}.$ (14)
The polynomials $Q_{n}$ are obtain by the usual recurrence relation:
$\epsilon Q_{n}=Q_{n+1}+\xi_{n}Q_{n-1}\quad\text{with}\quad\xi_{n}\equiv\frac{||Q_{n}||^{2}}{||Q_{n-1}||^{2}}$ (15)
We choose $Q_{0}=1$ and $Q_{1}=\epsilon$ as the initial condition. Note that even and odd polynomials are respectively symmetric and antisymmetric about the Fermi surface: $Q_{n}(-\epsilon)=(-1)^{n}Q_{n}(\epsilon)$. By contrast to this exact approach, the relaxation time approximations [19, 1] truncate the expansion in Eq. (13) to $n=0$, thereby neglecting the energy dependence of the quasiparticle distribution. The difference between the Abrikosov-Khalatnikov [19] and the variational formulation [1, 35, 29] lies in the treatment of the remaining $\epsilon$ dependence of the collision kernel: Abrikosov and Khalatnikov replaced it by its value in $\epsilon=0$ (in particular they approximate the quasiparticle lifetime by its value at the Fermi level $\Gamma(\epsilon)\approx\Gamma(0)$), while the variational formulation averages it over $\epsilon$.
Note that the polynomials $Q_{n}$ used here differ from the orthogonal polynomials of the momentum $p$ used at nonvanishing temperature [29, 20], and that the replacement Eq. (6) converts even/odd powers of $p^{2}$ into even/odd powers of $\epsilon$.
@@
$\Gamma_{nn^{\prime}}=\int_{-\infty}^{+\infty}\textrm{d}\epsilon g(\epsilon)\Gamma(\epsilon)\frac{Q_{n}(\epsilon)}{||Q_{n}||^{2}}Q_{n^{\prime}}(\epsilon)$ (16)
@@
$\!\!\!\mathcal{N}_{nn^{\prime}}^{l}\!=\!\!\int_{-\infty}^{+\infty}\!\!\!\!\!\!\!\!\textrm{d}\epsilon^{\prime}\textrm{d}\epsilon g(\epsilon)\int_{-1}^{1}\!\!\!\!\textrm{d}uP_{l}(u)\frac{Q_{n}(\epsilon)}{||Q_{n}||^{2}}\mathcal{N}_{\rm od}(\epsilon,\epsilon^{\prime},u)Q_{n^{\prime}}(\epsilon^{\prime})\\ =\frac{2}{2l+1}\mathcal{S}_{nn^{\prime}}\left((-1)^{l+n^{\prime}}-2\right)$ (17)
@@
$\mathcal{S}_{nn^{\prime}}=\int_{-\infty}^{+\infty}\textrm{d}\epsilon\textrm{d}\epsilon^{\prime}g(\epsilon)\mathcal{S}(\epsilon,\epsilon^{\prime})\frac{Q_{n}(\epsilon)}{||Q_{n}||^{2}}Q_{n^{\prime}}(\epsilon^{\prime}).$ (18)
Note that the subspaces of symmetric and antisymmetric functions of $\epsilon$ (even and odd $n$ respectively) are decoupled: $\Gamma_{nn^{\prime}}=\mathcal{N}_{nn^{\prime}}^{l}=0$ if ${n+n^{\prime}}$ is odd [42]. This is specific to the low-temperature limit where both the density of state $g(\epsilon)$ and energy integration domain are symmetric about the Fermi surface. Treating separately the odd and even orders, $\Gamma$ and $\mathcal{S}$ are tridiagonal matrices and can be expressed analytically as
@@
with $\xi_{n}=\frac{\pi^{2}n^{4}}{(2n+1)(2n-1)}$.
@@
$c\nu_{n}^{l}-\left[\frac{l+1}{2l+3}\nu_{n}^{l+1}+\frac{l}{2l-1}\nu_{n}^{l-1}\right]-\frac{2k_{\textrm{F}}a}{\pi}\delta_{l,1}\delta_{n,0}\nu_{0}^{0}\\ +\frac{\textrm{i}}{\pi\omega_{0}\tau}\sum_{n^{\prime}}\mathcal{M}_{nn^{\prime}}^{l}\nu_{n^{\prime}}^{l}=-\delta_{l,1}\delta_{n,0}$ (19)
@@
The conservation of the number of quasiparticles, energy, and momentum in a collision [see Eq. (4)] generates zero-energy eigenfunctions of the collision kernel. In our orthogonal basis, this simply translates into $\mathcal{M}_{0n^{\prime}}^{0}=\mathcal{M}_{n0}^{0}=0$, $\mathcal{M}_{1n^{\prime}}^{0}=\mathcal{M}_{n1}^{0}=0$ and $\mathcal{M}_{0n^{\prime}}^{1}=\mathcal{M}_{n0}^{1}=0$, respectively, The corresponding equations of motion on the conserved quantities $\nu_{0}^{0}$, $\nu_{0}^{1}$ and $\nu_{1}^{0}$ are
@@
$c\nu_{1}^{0}-\frac{\nu_{1}^{1}}{3}$ $=$ $0$ (23)
Physically $\nu_{0}^{0}=-{2\pi^{2}}\chi_{\rho}/mk_{\textrm{F}}$ is proportional to the density response $\chi_{\rho}=\delta\rho/U$, $\nu_{0}^{1}=-k_{\textrm{F}}\chi_{v}/2$ is proportional to the response of the velocity $\chi_{v}=v_{\parallel}/U$ with $\rho_{\rm eq}v_{\parallel}=({1}/{V})\sum_{\textbf{p}}(\textbf{p}\cdot\textbf{q}/mq)\delta n(\textbf{p})$, and $\nu_{1}^{0}=-{6}\chi_{e}/{mTk_{\textrm{F}}}$ is proportional to the response of the energy density $\chi_{e}=\delta e/U$ with $\delta e=(1/V)\sum_{\textbf{p}}(\textbf{p}^{2}/2m-\mu)\delta n(\textbf{p})$.
In the hydrodynamic limit, only these conserved quantities remain of order unity, while all the other components pick up one or several factors $\omega_{0}\tau$. To evaluate the power of a given component $\nu_{n}^{l}$ in $\omega_{0}\tau$, one should count the number of transport equations needed to reach a conserved quantity using the couplings appearing in Eq. (19), that is, $\nu_{n}^{l}\to\nu_{n}^{l\pm 1},\nu_{n^{\prime}}^{l}$ (with $n^{\prime}$ having the same parity as $n$). The components $\nu_{n}^{2}$ with $n$ even and $\nu_{n}^{1}$ with $n$ odd, which are directly coupled to the conserved quantities $\nu_{0}^{1}$ and $\nu_{1}^{0}$ respectively, are of order $O(\omega_{0}\tau)^{1}$, and the other components are subleading, as depicted by Fig. 1. In particular, the large $l$ components decay exponentially as $O(\omega_{0}\tau)^{l-1}$ or $O(\omega_{0}\tau)^{l}$, depending on the parity of $n$.
@@
The drive on the right-hand-side of Eq. (19) is coupled to $\nu_{0}^{1}$. The perturbation it generates is therefore symmetric in $\epsilon$, that is, $\nu_{n}^{l}=0$ for all $n$ odd. In particular the sound wave does not generate fluctuations of the energy density: $\nu_{1}^{0}=0$. To leading order in $\omega_{0}\tau$ the system Eqs. (21)–(22) describes an ideal hydrodynamic behavior, that is, an undamped resonance at $\omega_{q}=c_{1}q$ with the first sound velocity
@@
To study how the resonance deviates from $c_{1}$, one must compute the set $\vec{\nu}^{2}=(\nu_{n}^{2})_{n}$ of the non-conserved components in the $l=2$ subspace. Keeping the leading and subleading terms in this equation, we obtain
$\left(c+\frac{\textrm{i}}{\pi\omega_{0}\tau}\mathcal{M}^{2}\right)\vec{\nu}^{2}=\frac{2}{3}\nu_{0}^{1}\vec{u}_{0}+O(\omega_{0}\tau)^{2}$ (25)
where we have introduced the matrix $\mathcal{M}^{2}=(\mathcal{M}_{nn^{\prime}}^{l=2})_{nn^{\prime}}$ and the unit vectors $(\vec{u}_{n})_{n^{\prime}}=\delta_{nn^{\prime}}$. We have neglected in Eq. (25) the vector $\vec{\nu}^{3}$ and the components of $\vec{\nu}^{1}$ orthogonal to the zero eigenvector $\nu_{0}^{1}\vec{u}_{0}$: both are of order $O(\omega_{0}\tau)^{2}$ (see Fig. 1), and hence negligible compared to the subleading term $c\vec{\nu}^{2}$ in Eq. (25). As is noted in Ref. [42], the expansion in powers of $\omega_{0}\tau$ becomes very tedious as soon as higher Legendre components have to be taken into account. Fortunately this is not the case in our calculation to order $O(\omega_{0}\tau)^{2}$: the transport equation can be truncated to $l=2$, and only two collision times are needed to express $\nu_{0}^{2}$ from Eq. (25):
$\nu_{0}^{2}=-\frac{2\textrm{i}\omega_{0}\pi}{3}\nu_{0}^{1}\left[\tau_{\eta}+\textrm{i}\omega_{0}\pi ct_{\eta}^{2}\right]+O(\omega_{0}\tau)^{3}$ (26)
where we introduce the viscous collision time $\tau_{\eta}/\tau=\vec{u}_{0}\frac{1}{\mathcal{M}^{2}}\vec{u}_{0}$ and the second-order viscous time $t_{\eta}/\tau=\sqrt{\vec{u}_{0}\frac{1}{(\mathcal{M}^{2})^{2}}\vec{u}_{0}}$. These two parameters characterize the exact equations of motion of the density and parallel velocity within second-order hydrodynamics; a relaxation time approximation would not distinguish them since it amounts to replacing $\mathcal{M}^{2}$ by a number. In an exact calculation, they can be expressed as infinite series [13, 12] or as continued fractions [43], which converge rather slowly when truncated at $n=n^{\prime}=n_{\rm max}$: $\tau_{\eta}-\tau_{\eta}^{(n_{\rm max})}=O(1/n_{\rm max}^{2})$. Numerically, we obtain
$\tau_{\eta}\simeq 1.079\tau_{\eta}^{(0)}\simeq 0.102\tau\text{\ \ and\ \ }t_{\eta}\simeq 1.098\tau_{\eta}^{(0)}\simeq 0.104\tau$ (27)
where we have introduced the viscous collision time truncated at $n=0$: $\tau_{\eta}^{(0)}/\tau=1/\mathcal{M}^{l=2}_{00}=15/16\pi^{2}$.
@@
$\chi_{\rho}=\frac{mk_{\textrm{F}}}{6\pi^{2}}\frac{1}{c^{2}-\bar{c}_{1}^{2}+\frac{4\textrm{i}\omega_{0}\pi c}{15}\left(\tau_{\eta}+\textrm{i}\omega_{0}\pi ct_{\eta}^{2}\right)}.$ (28)
The pole $z_{1}$ of this Lorentzian is located at:
$z_{1}=\bar{c}_{1}-\frac{2\textrm{i}\pi}{15}\omega_{0}\tau_{\eta}+\frac{2\pi^{2}\omega_{0}^{2}}{225\bar{c}_{1}}(15\bar{c}_{1}^{2}t_{\eta}^{2}-\tau_{\eta}^{2})$ (29)
@@
$\omega_{q}\!=\!c_{1}q\left[1+\theta(\bar{c}_{1})\epsilon_{\textrm{F}}^{2}\tau^{2}\left(\frac{q}{mc_{1}}\right)^{2}\!\!\!+O(\omega_{0}\tau)^{3}\right]$ (30)
with $\theta(c)\simeq 0.057c^{2}-0.0037$. The first deviation from the linear spectrum $c_{1}q$ is thus proportional to $q^{3}$. The dispersion is positive at weak coupling since $\theta(\bar{c}_{1})\simeq 0.0153$ for $\bar{c}_{1}=1/\sqrt{3}$. We note that $\theta(c)$ is negative for $c<c_{\rm inv}\simeq 0.25$, such that an inversion of the sign of the dispersion may occur in settings where the ratio $c_{1}/v_{\textrm{F}}$ is lower than $c_{\rm inv}$.
@@
$\Gamma_{q}=\frac{2}{3m\rho}\eta q^{2}\text{ with }\eta=\frac{2\pi}{5}\rho\epsilon_{\textrm{F}}\tau_{\eta}\simeq 0.129\rho\epsilon_{\textrm{F}}\tau.$ (31)
@@
In passing, the equation on the energy density Eq. (23) allows us to compute the thermal diffusivity of the gas. By inverting the matrix $\mathcal{M}^{1}$ in the subspace of the $\nu_{n}^{1}$ with $n$ odd, where it has no conserved quantities, we derive
$\nu_{1}^{1}=-\textrm{i}\pi\omega_{0}\tau_{\kappa}\nu_{1}^{0}\implies\left(\omega+\textrm{i}{D_{\kappa}}q^{2}\right)\nu_{1}^{0}=0\\ \text{ with }D_{\kappa}=\frac{\pi}{3}{v_{\textrm{F}}^{2}}\tau_{\kappa}\simeq 0.0623{v_{\textrm{F}}^{2}\tau}.$ (32)
We have introduced the thermal diffusion time $\tau_{\kappa}/\tau=\vec{u}_{1}\frac{1}{\mathcal{M}^{1}}\vec{u}_{1}$. The relaxation time approximation of this time $\tau_{\kappa}^{(0)}/\tau=15/32\pi^{2}$ is much worse than for the viscous time $\tau_{\eta}$, with the error reaching $25\%$.
Finally, we note that the nonvanishing isotropic components of the collision kernel $\mathcal{M}^{0,\perp}=(\mathcal{M}^{0})_{n>1,n^{\prime}>1}$, which are usually associated with the bulk viscosity $\zeta$, do not affect the propagation of sound, even within second-order hydrodynamics. This means the bulk viscosity appears in the hydrodynamic equations of a Fermi liquid. Nevertheless, the comparison of the exact to approximate matrix elements $(1/\mathcal{M}^{0,\perp})_{22}\simeq 1.41/\mathcal{M}^{0}_{22}$ suggests that the relaxation time approximations estimate the bulk viscosity quite poorly.

### v3

@@ III Hydrodynamic limit
In this regime, collisions bring the quasiparticle distribution back to equilibrium much faster than the typical time $1/\omega_{0}$ at which the sound wave evolves. Only the few components of the distribution that are not affected by collisions (i.e. those that belong to the zero-energy space of the collision kernel) remain significantly excited. Eq. (13) defines a dynamical regime, accessible from any equilibrium state of the phase diagram (i.e. for any value of $a$ and $T$). Here we have sent $a$ and $T$ to 0 at fixed $\omega_{0}\tau$ and finally expanded the transport equation in powers of $\omega_{0}\tau$. To guarantee that the system remains in the hydrodynamic regime, it is required to set $q/k_{\rm F}=2(T/T_{\rm F})^{2}(k_{\rm F}a)^{2}\omega_{0}\tau_{\sigma}$ in such a way that $\omega_{0}\tau_{\sigma}\ll 1$. For a finite-size system (of size $L$), $q$ is limited to $q=\pi/L$ which in typical experiments is $\approx 0.01k_{\rm F}$.
The usual expansion to order $O(\omega_{0}\tau)^{1}$ yields the transport coefficients entering the Fermi-liquid version of the Navier-Stokes equations. Here, we will push to second-order in the hydrodynamic expansion, that is $O(\omega_{0}\tau)^{2}$, and retrieve the sound dispersion from the quantum analogue of Burnett’s hydrodynamics [43, 44].
To obtain the exact solution of the transport equation in the hydrodynamic limit, we expand the distribution $\nu$ over a basis of orthogonal polynomials
$\nu(\epsilon,\theta)=\sum_{n,l=0}^{+\infty}\nu_{n}^{l}P_{l}(\cos\theta)Q_{n}(\epsilon)$ (14)
@@
[…]
The polynomials $Q_{n}$ are obtained by the usual recurrence relation:
[…]
We choose $Q_{0}=1$ and $Q_{1}=\epsilon$ as the initial condition111 Note that polynomials $Q_{n}$ used here differ from the orthogonal polynomials of the momentum $p$ used at nonvanishing temperature [30, 21]. The replacement Eq. (6) converts even/odd powers of $p^{2}$ into even/odd powers of $\epsilon$.. Note that even and odd polynomials are respectively symmetric and antisymmetric about the Fermi surface: $Q_{n}(-\epsilon)=(-1)^{n}Q_{n}(\epsilon)$.
By contrast to our exact approach, the relaxation time approximations [20, 1] truncate the expansion in Eq. (14) to $n=0$, thereby neglecting the energy dependence of the quasiparticle distribution. The difference between the Abrikosov-Khalatnikov [20] and the variational formulation [1, 38, 30] lies in the treatment of the remaining $\epsilon$ dependence of the collision kernel: Abrikosov and Khalatnikov replaced it by its value in $\epsilon=0$ (in particular they approximate the quasiparticle lifetime by its value at the Fermi level $\Gamma(\epsilon)\approx\Gamma(0)$), while the variational formulation averages it over $\epsilon$.
@@
[…]
@@
[…]
@@
[…]
Note that the subspaces of symmetric and antisymmetric functions of $\epsilon$ (even and odd $n$ respectively) are decoupled: $\Gamma_{nn^{\prime}}=\mathcal{N}_{nn^{\prime}}^{l}=0$ if $n+n^{\prime}$} is odd [48]. This property is specific to the low-temperature limit where both the density of state $g(\epsilon)$ and energy integration domain are symmetric about the Fermi surface. Treating separately the odd and even orders, $\Gamma$ and $\mathcal{S}$ are tridiagonal matrices and can be expressed analytically as
@@
with $\xi_{n}=\frac{\pi^{2}n^{4}}{(2n+1)(2n-1)}$ as can be shown recursively.
@@
$c\nu_{n}^{l}-\left[\frac{l+1}{2l+3}\nu_{n}^{l+1}+\frac{l}{2l-1}\nu_{n}^{l-1}\right]-\frac{2k_{\textrm{F}}a}{\pi}\delta_{l,1}\delta_{n,0}\nu_{0}^{0}\\ +\frac{\textrm{i}}{\omega_{0}\tau}\sum_{n^{\prime}}\mathcal{M}_{nn^{\prime}}^{l}\nu_{n^{\prime}}^{l}=-\delta_{l,1}\delta_{n,0}$ (21)
@@
The conservation of the number of quasiparticles, energy, and momentum in a collision [see Eq. (4)] generates zero-energy eigenfunctions of the collision kernel. In our orthogonal basis, this translates respectively into $\mathcal{M}_{0n^{\prime}}^{0}=\mathcal{M}_{n0}^{0}=0$, $\mathcal{M}_{1n^{\prime}}^{0}=\mathcal{M}_{n1}^{0}=0$ and $\mathcal{M}_{0n^{\prime}}^{1}=\mathcal{M}_{n0}^{1}=0$ for all $n,n^{\prime}\in\mathbb{N}$. The corresponding equations of motion on the conserved quantities $\nu_{0}^{0}$, $\nu_{0}^{1}$ and $\nu_{1}^{0}$ are
@@
$c\nu_{1}^{0}-\frac{\nu_{1}^{1}}{3}$ $=$ $0.$ (25)
To give a physical interpretation to the three conserved components of the distribution $\nu$, we relate them to the fluctuations of the density $\delta\rho$, longitudinal velocity $v_{\parallel}$ and energy density $e$ of the gas:
$\delta\rho(\textbf{q},\omega)$ $=$ $\frac{1}{V}\sum_{\textbf{p}}\delta n(\textbf{p},\textbf{q},\omega)$ (26)
$\rho_{\rm eq}v_{\parallel}(\textbf{q},\omega)$ $=$ $\frac{1}{V}\sum_{\textbf{p}}\frac{\textbf{p}\cdot\textbf{q}}{mq}\delta n(\textbf{p},\textbf{q},\omega)$ (27)
$\delta e(\textbf{q},\omega)$ $=$ $\frac{1}{V}\sum_{\textbf{p}}\left(\frac{\textbf{p}^{2}}{2m}-\mu\right)\delta n(\textbf{p},\textbf{q},\omega)$ (28)
Remarking that $\nu$ is scaled to the intensity $U$ of the drive (Eq. (6)), one expresses the conserved components in terms of the linear responses $\chi$ of the system:
$\nu_{0}^{0}$ $=$ $-\frac{2\pi^{2}}{mk_{\textrm{F}}}{\chi_{\rho}}\text{, }\qquad\chi_{\rho}\equiv\frac{\delta\rho}{U}$ (29)
$\nu_{0}^{1}$ $=$ $-2k_{\textrm{F}}\chi_{v}\text{, }\qquad\chi_{v}\equiv\frac{v_{\parallel}}{U}$ (30)
$\nu_{1}^{0}$ $=$ $-\frac{6}{mTk_{\textrm{F}}}{\chi_{e}}\text{, }\qquad\chi_{e}\equiv\frac{\delta e}{U}$ (31)
In the hydrodynamic limit, only these conserved quantities remain of order unity, while all the other components pick up one or several factors $\omega_{0}\tau$. To evaluate the power of a given component $\nu_{n}^{l}$ in $\omega_{0}\tau$, one should count the number of projected transport equations (21) needed to reach a conserved quantity using the couplings appearing in Eq. (21), that is, $\nu_{n}^{l}\to\nu_{n}^{l\pm 1}$ and, via the collision kernel, $\nu_{n}^{l}\to\nu_{n^{\prime}}^{l}$ (with $n^{\prime}$ having the same parity as $n$). The components $\nu_{n}^{2}$ with $n$ even and $\nu_{n}^{1}$ with $n$ odd, which are directly coupled to the conserved quantities $\nu_{0}^{1}$ and $\nu_{1}^{0}$ respectively, are of order $O(\omega_{0}\tau)^{1}$. All the other components are subleading, as depicted by Fig. 1. In particular, the large $l$ components decay exponentially as $O(\omega_{0}\tau)^{l-1}$ or $O(\omega_{0}\tau)^{l}$, depending on the parity of $n$.
@@
IV First-sound waves
The drive on the right-hand-side of Eq. (21) is coupled to $\nu_{0}^{1}$. The perturbation it generates is therefore symmetric in $\epsilon$, that is, $\nu_{n}^{l}=0$ for all $n$ odd. In particular the sound wave does not generate fluctuations of the energy density: $\nu_{1}^{0}=0$. To leading order in $\omega_{0}\tau$ the system of Eqs. (23)–(24) describes an ideal hydrodynamic behavior, that is, an undamped resonance at $\omega_{q}=c_{1}q$ with the first sound velocity
@@ IV First-sound waves
To study how the resonance deviates from $c_{1}$, one must compute the set $\vec{\nu}^{2}=(\nu_{n}^{2})_{n\in\mathbb{N}}$ of the non-conserved components in the $l=2$ subspace, that we write as an infinite-dimension vector. Keeping the leading and subleading terms in Equation (21) for $l=2$, we obtain
$\left(c+\frac{\textrm{i}}{\omega_{0}\tau}\mathcal{M}^{2}\right)\vec{\nu}^{2}=\frac{2}{3}\nu_{0}^{1}\vec{u}_{0}+O(\omega_{0}\tau)^{2}$ (33)
where we have introduced the matrix $\mathcal{M}^{2}=(\mathcal{M}_{nn^{\prime}}^{l=2})_{n,n^{\prime}\in\mathbb{N}}$ and the unit vectors $(\vec{u}_{n})_{n^{\prime}\in\mathbb{N}}=(\delta_{nn^{\prime}})_{n^{\prime}}$. We have neglected in Eq. (33) the vector $\vec{\nu}^{3}$ and the components of $\vec{\nu}^{1}$ orthogonal to the zero eigenvector $\nu_{0}^{1}\vec{u}_{0}$: both are of order $O(\omega_{0}\tau)^{2}$ (see Fig. 1), and hence negligible compared to the subleading term $c\vec{\nu}^{2}$ (of order $O(\omega_{0}\tau)$) in Eq. (33). As is noted in Ref. [48], the expansion in powers of $\omega_{0}\tau$ becomes very tedious as soon as higher Legendre components have to be taken into account. Fortunately this is not the case in our calculation to order $O(\omega_{0}\tau)^{2}$: the transport equation can be truncated to $l=2$, and only two collision times are needed to express $\nu_{0}^{2}$ from Eq. (33):
$\nu_{0}^{2}=-\frac{2\textrm{i}\omega_{0}}{3}\nu_{0}^{1}\left[\tau_{\eta}+\textrm{i}\omega_{0}ct_{\eta}^{2}\right]+O(\omega_{0}\tau)^{3}$ (34)
where we introduce the viscous collision time $\tau_{\eta}/\tau=\vec{u}_{0}\frac{1}{\mathcal{M}^{2}}\vec{u}_{0}$ and the second-order viscous time $t_{\eta}/\tau=\sqrt{\vec{u}_{0}\frac{1}{(\mathcal{M}^{2})^{2}}\vec{u}_{0}}$. These two parameters characterize the Burnett hydrodynamical equations on the density and parallel velocity; a relaxation time approximation would not distinguish them since it amounts to replacing $\mathcal{M}^{2}$ by a number. In previous exact calculations, $\tau_{\eta}$ has been expressed as an infinite serie [14, 13] or as continued fraction [49], which converge rather slowly when truncated at $n=n^{\prime}=n_{\rm max}$: $\tau_{\eta}-\tau_{\eta}^{(n_{\rm max})}=O(1/n_{\rm max}^{2})$. Numerically, we obtain
$\tau_{\eta}\simeq 1.079\tau_{\eta}^{(0)}\simeq 1.288\tau_{\sigma}\text{\ \ and\ \ }t_{\eta}\simeq 1.098\tau_{\eta}^{(0)}\simeq 1.311\tau_{\sigma}$ (35)
where we have introduced the viscous collision time truncated at $n=0$: $\tau_{\eta}^{(0)}/\tau_{\sigma}=4\pi/\mathcal{M}^{l=2}_{00}=15/4\pi$.
@@
$\chi_{\rho}=\frac{mk_{\textrm{F}}}{6\pi^{2}}\frac{1}{c^{2}-\bar{c}_{1}^{2}+\frac{4\textrm{i}\omega_{0}c}{15}\left(\tau_{\eta}+\textrm{i}\omega_{0}ct_{\eta}^{2}\right)}.$ (36)
The pole $z_{1}$ of $\chi_{\rho}(c)$ is located at:
$z_{1}=\bar{c}_{1}-\frac{2\textrm{i}}{15}\omega_{0}\tau_{\eta}+\frac{2\omega_{0}^{2}}{225\bar{c}_{1}}(15\bar{c}_{1}^{2}t_{\eta}^{2}-\tau_{\eta}^{2})$ (37)
@@
$\omega_{q}\!=\!c_{1}q\left[1+\theta(\bar{c}_{1})\epsilon_{\textrm{F}}^{2}\tau_{\sigma}^{2}\left(\frac{q}{mc_{1}}\right)^{2}\!\!\!+O(\omega_{0}\tau)^{3}\right]$ (38)
with $\theta(c)\simeq 0.917c^{2}-0.059$. The first deviation from the linear spectrum $c_{1}q$ is thus proportional to $q^{3}$. The dispersion is positive at weak coupling since $\theta(\bar{c}_{1})\simeq 0.247$ for $\bar{c}_{1}=1/\sqrt{3}$. The function $\theta$ vanishes and changes sign at $c_{\rm inv}\simeq 0.25$ such that an inversion of the sign of the dispersion may occur in settings where the ratio $c_{1}/v_{\textrm{F}}$ is lower than $c_{\rm inv}$.
@@
$\Gamma_{q}=\frac{2}{3m\rho}\eta q^{2}\text{ with }\eta=\frac{2}{5}\rho\epsilon_{\textrm{F}}\tau_{\eta}\simeq 0.5153\rho\epsilon_{\textrm{F}}\tau_{\sigma}.$ (39)
@@
V Thermal conductivity
Equation (25) on the energy density allows us to compute the thermal conductivity of the gas. At low temperature, the energy current $\nu_{1}^{1}$ exists only if the quasiparticle distribution has components odd in $n$, that is if there is an energy-imbalance across the Fermi surface, which is not the case in a sound wave. We compute this energy current by writing down the transport equation for $n$ odd and $l=1$:
$c\nu_{n}^{1}+\frac{\textrm{i}}{\omega_{0}\tau}\sum_{n^{\prime}}\mathcal{M}_{nn^{\prime}}^{1}\nu_{n^{\prime}}^{1}-\nu_{n}^{0}-\frac{2}{5}\nu_{n}^{2}=0$ (40)
To leading order in $\omega_{0}\tau$, this gives
$\frac{\textrm{i}}{\omega_{0}\tau}\mathcal{M}^{1}\vec{\nu}_{\rm odd}^{1}=\nu_{1}^{0}\vec{u}_{1}+O(\omega_{0}\tau)$ (41)
where $\vec{\nu}_{\rm odd}^{1}=(0,\nu_{1}^{1},0,\nu_{3}^{1},0,\ldots)$ is the odd component of $\vec{\nu}^{1}$, and we recall that $\mathcal{M}$ does not couple the even and odd components. We solve Eq. (41) for $\nu_{1}^{1}$, which yields
$\nu_{1}^{1}=-\textrm{i}\omega_{0}\tau_{\kappa}\nu_{1}^{0}\quad\text{with}\quad\frac{\tau_{\kappa}}{\tau_{\sigma}}=\vec{u}_{1}\frac{4\pi}{\mathcal{M}^{1}}\vec{u}_{1}\simeq 0.7478$ (42)
Going back to Eq. (25) on the energy density, we obtain the heat diffusion equation
$\left(\omega+\textrm{i}{D_{\kappa}}q^{2}\right)\nu_{1}^{0}=0\text{ with }D_{\kappa}=\frac{1}{3}{v_{\textrm{F}}^{2}}\tau_{\kappa}\simeq 0.249{v_{\textrm{F}}^{2}\tau_{\sigma}}.$ (43)
The thermal conductivity $\kappa$ is related to $D_{\kappa}$ by $\kappa=c_{V}D_{\kappa}$ with $c_{V}=mp_{\rm F}T/3$ the heat capacity. The relaxation time approximation $\tau_{\kappa}^{(0)}/\tau_{\sigma}=15/8\pi$ turns out much worse for the thermal conductivity than for the viscosity, with the error reaching $25\%$. We are thus in neat disagreement with the claim of Ref. [30] that $|\kappa-\kappa_{(0)}|/\kappa<2\%$.
VI Conclusion
Dispersion of sound was extensively studied in classical gases, where it is very well captured by second-order (Burnett) hydrodynamics [50]. We have shown here that a gas of fermions obeys a very similar kind of Burnett hydrodynamics, although the presence of a Fermi sea leads to a $1/T^{2}$ behavior of the collision time, as opposed to $\tau\propto\sqrt{T}$ in a classical gas. A quantum gas is also affected by a quantum Vlasov force on the left-hand side of the transport equation. To leading order in $a$, this force is the gradient of $g\rho(\textbf{r})$, and therefore independent of the quasiparticle momentum p. This may not be true to higher order in the interaction strength, and the effect on hydrodynamic transport equations remains to be investigated.
Experiments on ultracold gases have often measured transport coefficients [23, 51, 52, 53, 54, 55], and thus seem mature to investigate departures from standard Navier-Stokes hydrodynamics. The dispersion of first sound could be measured by tracking the variations of the resonance frequency in the density-density response function with the excitation wavenumber $q$. Alternatively, this resonance frequency could be compared to the speed of sound known from the equation of state. Dispersive effects also affect the real time dynamics, such as the equilibration after a quench [56], or the propagation of shock waves [42, 57, 58].